OpenStudy (anonymous):

The distance between two cities is ninety miles, and a woman drives from one city to the other at a rate of 45 mph. At what rate must she return if the total travel time is three hours and forty minutes? 27 mph 45 mph 54 mph

OpenStudy (anonymous):

54

sam (.sam.):

distance = 90 mi ; v = 45 mph ; time = d/v = 2 hrs. distance = 90 mi ; time = [3 2/3 - 2] = 1 2/3 hrs v = 90/(5/3) = 54 mph

OpenStudy (anonymous):

you can use this formula: rate = distance/time